OpenStudy (anonymous):

Find all values of x for which the series converges, and find the sum of the series for those values of x.

OpenStudy (anonymous):

1/(x^2)+4/(x^3)+16/(x^4)+64/(x^5)+256/(x^6)

OpenStudy (anonymous):

converges for x < ? or x> ?

OpenStudy (perl):

what is the original series

OpenStudy (perl):

a series has infinite terms, does this series go on

OpenStudy (anonymous):

yes :)

OpenStudy (perl):

ok it looks like your series is $$ \Large{\sum_{n=0}^{\infty} \frac{4^{n}}{x^{n+2}} } $$

OpenStudy (perl):

to find where it converges , we can use ratio test

OpenStudy (perl):

alternatively, you can rewrite the series

OpenStudy (perl):

ok it looks like your series is $$ \Large{\sum_{n=0}^{\infty} \frac{4^{n}}{x^{n+2}} = \sum_{n=0}^{\infty} \frac{4^{n}}{x^{n}\cdot x^2}= \sum_{n=0}^{\infty} (\frac{4}{x})^n\cdot \frac{1}{x^2} \\ = } $$

OpenStudy (anonymous):

i got it ! thank you :)